Please provide clear work. Thanks!Consider the following IS-LM model in a closed economy: C = 335 + 0.3YD L = 130 + 0.15Y – 850i G = 350 T = 245 i= 0.04 M/P =
Answer #1

(= 335 to 3 Yo I = 130 +0.1SY - 8501 G - 350 T-2uS -- M 3.67 -9 250 c = Consumption Yo = V-T) disposable disposable Real GDP

HS - Investment hout Expenditure T Tax ek Money supply (ط) Money demand = 3.69-925o i Is equation derivation In Equilibrium Y

IS equation 1348.18 - Isus.usi Y